# iQUEST: An Iterative Question-Guided Framework for Knowledge Base Question Answering

### v3

@@ Abstract
While Large Language Models (LLMs) excel at many natural language processing tasks, they often suffer from factual inaccuracies in knowledge-intensive scenarios. Integrating external knowledge resources, particularly knowledge graphs (KGs), provides a transparent and updatable foundation for more reliable reasoning. Knowledge Base Question Answering (KBQA), which queries and reasons over KGs, is central to this effort, especially for complex, multi-hop queries. However, multi-hop reasoning poses two key challenges: (1) maintaining coherent reasoning paths, and (2) avoiding prematurely discarding critical multi-hop connections. To address these issues, we introduce iQUEST, a question-guided KBQA framework that iteratively decomposes complex queries into simpler sub-questions, ensuring a structured and focused reasoning trajectory. Additionally, we integrate a Graph Neural Network (GNN) to look ahead and incorporate 2-hop neighbor information at each reasoning step. This dual approach strengthens the reasoning process, enabling the model to explore viable paths more effectively. Detailed experiments demonstrate the consistent improvement delivered by iQUEST across four benchmark datasets and four LLMs.
iQUEST: An Iterative Question-Guided Framework for Knowledge Base Question Answering
@@ 1 Introduction
Large language models (LLMs) have achieved remarkable success across diverse Natural Language Processing (NLP) tasks, yet they often exhibit hallucinations and factual errors in specialized, knowledge-intensive domains Huang et al. (2024); Martino et al. (2023); Minaee et al. (2024). Fine-tuning LLMs on curated datasets can embed domain-specific knowledge into model parameters, but it is computationally expensive, difficult to update, and opaque in terms of interpretability Hu et al. (2023). As an alternative, Retrieval-Augmented Generation (RAG) techniques query external resources during inference, thereby reducing reliance on repeated retraining Gao et al. (2023). Among these resources, knowledge graphs (KGs) offer structured and trustworthy information that is directly verifiable and maintainable, an essential feature for high-stakes scenarios such as healthcare and autonomous driving Wen et al. (2024).
@@
In this context, Knowledge Base Question Answering (KBQA) has emerged as a crucial research direction, aiming to query and reason over KGs to answer natural language questions. Complex KBQA tasks often require multi-hop reasoning or multiple constraints Lan et al. (2022), as illustrated by questions like “In the North Pacific region, what is the official flower of the affected area of the cyclone Tropical Storm Fabio?” Answering such queries necessitates iterative exploration of entities, relations, and constraints within large, heterogeneous KGs.
A straightforward way to handle complex queries is the Chain-of-Thought (CoT) Wei et al. (2022) framework, which makes each inference step explicit. Building on this idea, some methods treat LLMs as agents that iteratively explore the KG from a topic node (e.g., ToG Sun et al. (2024), Interactive-KBQA Xiong et al. (2024)), while others combine a smaller model with an LLM to gather evidence before synthesizing the final answer (e.g., KG-CoT Zhao et al. (2024)). Another direction focuses on sub-question decomposition, either by inserting delimiters into the original query Huang et al. (2023) or by using an LLM to break down the query for a fine-tuned smaller model Yixing et al. (2024). Although these approaches have shown promise, multi-step reasoning over KGs still face two key challenges.
@@
In complex problem-solving scenarios, people often lose focus if their attention is not guided continuously. However, research indicates that by posing and solving smaller questions, humans can maintain higher levels of attention during the reasoning process, thereby improving decision-making and problem-solving Salmon and Barrera (2021); Tofade et al. (2013). Inspired by this insight, we propose to guide LLMs in multi-step KBQA by iteratively posing simpler sub-questions, thereby maintaining a clear reasoning trajectory over multiple hops.
@@ 2.1 KBQA
Knowledge Base Question Answering (KBQA) aims to leverage a KG to generate answers. Traditional methods are generally IR-based or SP-based: the former retrieves candidate entities and relations from a KG and then ranks them Saxena et al. (2020); Dai et al. (2023); Chen et al. (2022), whereas the latter converts natural language questions into executable structured queries (e.g., SPARQL) Das et al. (2021); Ye et al. (2022).
Recently, LLMs have introduced new possibilities. Several approaches integrate LLMs in different ways: Interactive-KBQA Xiong et al. (2024) iteratively refines SPARQL queries; ToG Sun et al. (2024) uses an LLM agent to explore a KG for evidence; FlexKBQA Li et al. (2024) synthesizes training data for smaller models; KG-CoT Zhao et al. (2024) fuses small-model reasoning with LLM-based answer finalization. Other work uses tool-like interfaces for external KG reasoning Jiang et al. (2023) or multi-agent strategies Zong et al. (2024) to enhance KBQA performance.
@@ 2.2 Question Decomposition
A central challenge in complex KBQA is decomposing the query into tractable sub-questions. Early methods used neural architectures (e.g., Seq2Seq or BERT+LSTM) to split questions, but often lacked robustness or altered entity semantics Zhang et al. (2019); Gu and Su (2022). Later work introduced explicit delimiters to preserve meaning and sequentially answer sub-questions Huang et al. (2023). Increasingly, LLMs have been adopted for decomposition and subsequent reasoning Yixing et al. (2024). However, most focus on segmentation rather than adaptive inference. In contrast, our approach uses targeted intermediate prompts to guide and validate each reasoning step, harnessing the LLM’s capabilities more comprehensively.
@@ 2.3 Entity Exploration
Many KBQA systems rely on 1-hop neighbors for candidate retrieval, employing text similarity (e.g., BM25), direct semantic relevance, or MLP-based scoring Sun et al. (2024); Xiong et al. (2024); Zhao et al. (2024); Yixing et al. (2024). These methods often overlook 2-hop neighbors, potentially missing critical inference paths. Our work addresses this gap by explicitly incorporating 2-hop neighbors into the search process, enabling more thorough exploration and improved answer derivation.
@@ 4.2 Two-Hop Based Entity Exploration
Two-Hop Neighbor Aggregation with GNN. While 1-hop neighbors can be useful, they may not provide sufficient semantic context. To incorporate information from 2-hop neighbors, we perform another round of SPARQL queries to retrieve these nodes. We then employ a GNN to aggregate these 2-hop neighborhood signals, as shown in Step (3) in Algotithm 1. Specifically, we first convert textual features of each node (e.g., entity descriptions or relevant text) into dense vector representations using BERT. Let $\mathbf{h}_{1h}$ denote the representation for a 1-hop neighbor $e_{1h}$, while $\mathbf{h}_{2h}$ stands for a 2-hop neighbor $e_{2h}$. We adopt a GraphSAGE-like aggregation Hamilton et al. (2017) to update the representation of each 1-hop neighbor with its 2-hop neighbors:
@@ 5 Experiments
Question-Decomposion Huang et al. (2023) 72.8 - - -
Chain-of-Question Yixing et al. (2024) 78.8 78.10 - -
KG-CoT Zhao et al. (2024) 62.3 84.90 68.00 -
FlexKBQA Li et al. (2024) - 46.20 - 68.90
Interactive-KBQA Xiong et al. (2024) 49.07 71.2 - -
ToG Sun et al. (2024) 69.5 82.6 57.90 81.4
@@
We evaluate our approach on four standard KBQA datasets, all of which require multi-hop reasoning: ComplexWebQuestions (CWQ) Talmor and Berant (2018), WebQuestionsSP (WebQSP) Yih et al. (2016), WebQuestions Berant et al. (2013), and GrailQA Gu et al. (2021). Among them, GrailQA is also used to assess the generalization ability of the retrieval model. All these datasets are based on the Freebase knowledge graph Bollacker et al. (2008). Following previous studies, we adopt Hit@1 score as the primary evaluation metric.
Implementation Details. For the LLMs, we utilize GPT-4o (gpt-4o-2024-05-13), one of the most advanced general-purpose LLMs, alongside DeepSeek-R1 (70B), a recently released, powerful open-source reasoning model Guo et al. (2025). Additionally, we incorporate LLaMA 3.1-70B (hereafter referred to as LLaMA 70B) and a smaller-scale model, LLaMA 3.2-3B (referred to as LLaMA 3B), for comparative analysis.

### v4

# An Iterative Question-Guided Framework for Knowledge Base Question Answering

@@ Abstract
Large Language Models (LLMs) excel in many natural language processing tasks but often exhibit factual inconsistencies in knowledge-intensive settings. Integrating external knowledge resources, particularly knowledge graphs (KGs), provides a transparent and updatable foundation for more reliable reasoning. Knowledge Base Question Answering (KBQA), which queries and reasons over KGs, is central to this effort, especially for complex, multi-hop queries. However, multi-hop reasoning poses two key challenges: (1) maintaining coherent reasoning paths, and (2) avoiding prematurely discarding critical multi-hop connections. To tackle these challenges, we introduce iQUEST, a question-guided KBQA framework that iteratively decomposes complex queries into simpler sub-questions, ensuring a structured and focused reasoning trajectory. Additionally, we integrate a Graph Neural Network (GNN) to look ahead and incorporate 2-hop neighbor information at each reasoning step. This dual approach strengthens the reasoning process, enabling the model to explore viable paths more effectively. Detailed experiments demonstrate the consistent improvement delivered by iQUEST across four benchmark datasets and four LLMs.
An Iterative Question-Guided Framework for Knowledge Base Question Answering
@@ 1 Introduction
Large language models (LLMs) have achieved remarkable success across diverse Natural Language Processing (NLP) tasks, yet they often exhibit hallucinations and factual errors in specialized, knowledge-intensive domains huang2024survey; martino2023knowledge; minaee2024large. Fine-tuning LLMs on curated datasets can embed domain-specific knowledge into model parameters, but it is computationally expensive, difficult to update, and opaque in terms of interpretability hu2023llm. As an alternative, Retrieval-Augmented Generation (RAG) techniques query external resources during inference, thereby reducing reliance on repeated retraining gao2023retrieval. Among these resources, knowledge graphs (KGs) offer structured and trustworthy information that is directly verifiable and maintainable, an essential feature for high-stakes scenarios such as healthcare and autonomous driving wen-etal-2024-mindmap.
@@
In this context, Knowledge Base Question Answering (KBQA) has emerged as a crucial research direction, aiming to query and reason over KGs to answer natural language questions. Complex KBQA tasks often require multi-hop reasoning or multiple constraints lan2022complex, as illustrated by questions like “In the North Pacific region, what is the official flower of the affected area of the cyclone Tropical Storm Fabio?” Answering such queries necessitates iterative exploration of entities, relations, and constraints within large, heterogeneous KGs.
A straightforward way to handle complex queries is the Chain-of-Thought (CoT) wei2022chain framework, which makes each inference step explicit. Building on this idea, some methods treat LLMs as agents that iteratively explore the KG from a topic node (e.g., ToG sun2024thinkongraph, Interactive-KBQA xiong-etal-2024-interactive), while others combine a smaller model with an LLM to gather evidence before synthesizing the final answer (e.g., KG-CoT zhao2024kg). Another direction focuses on sub-question decomposition, either by inserting delimiters into the original query huang2023question or by using an LLM to break down the query for a fine-tuned smaller model yixing2024chain. Although these approaches have shown promise, multi-step reasoning over KGs still face two key challenges.
@@
In complex problem-solving scenarios, people often lose focus if their attention is not guided continuously. However, research indicates that by posing and solving smaller questions, humans can maintain higher levels of attention during the reasoning process, thereby improving decision-making and problem-solving salmon2021intentional; tofade2013best. Inspired by this insight, we propose to guide LLMs in multi-step KBQA by iteratively posing simpler sub-questions, thereby maintaining a clear reasoning trajectory over multiple hops.
@@ 2.1 KBQA
Knowledge Base Question Answering (KBQA) aims to leverage a KG to generate answers. Traditional methods are generally IR-based or SP-based: the former retrieves candidate entities and relations from a KG and then ranks them saxena2020improving; dai2023fedgamma; chen2022temporal, whereas the latter converts natural language questions into executable structured queries (e.g., SPARQL) das2021case; ye2022rng.
Recently, LLMs have introduced new possibilities. Several approaches integrate LLMs in different ways: Interactive-KBQA xiong-etal-2024-interactive iteratively refines SPARQL queries; ToG sun2024thinkongraph uses an LLM agent to explore a KG for evidence; FlexKBQA li2024flexkbqa synthesizes training data for smaller models; KG-CoT zhao2024kg fuses small-model reasoning with LLM-based answer finalization. Other work uses tool-like interfaces for external KG reasoning jiang2023structgpt or multi-agent strategies zong-etal-2024-triad to enhance KBQA performance.
@@ 2.2 Question Decomposition
A central challenge in complex KBQA is decomposing the query into tractable sub-questions. Early methods used neural architectures (e.g., Seq2Seq or BERT+LSTM) to split questions, but often lacked robustness or altered entity semantics zhang2019complex; gu-su-2022-arcaneqa. Later work introduced explicit delimiters to preserve meaning and sequentially answer sub-questions huang2023question. Increasingly, LLMs have been adopted for decomposition and subsequent reasoning yixing2024chain. However, most focus on segmentation rather than adaptive inference. In contrast, our approach uses targeted intermediate prompts to guide and validate each reasoning step, harnessing the LLM’s capabilities more comprehensively.
@@ 2.3 Entity Exploration
Many KBQA systems rely on 1-hop neighbors for candidate retrieval, employing text similarity (e.g., BM25), direct semantic relevance, or MLP-based scoring sun2024thinkongraph; xiong-etal-2024-interactive; zhao2024kg; yixing2024chain. These methods often overlook 2-hop neighbors, potentially missing critical inference paths. Our work addresses this gap by explicitly incorporating 2-hop neighbors into the search process, enabling more thorough exploration and improved answer derivation.
@@ 4.2 Two-Hop Based Entity Exploration
Two-Hop Neighbor Aggregation with GNN. While 1-hop neighbors can be useful, they may not provide sufficient semantic context. To incorporate information from 2-hop neighbors, we perform another round of SPARQL queries to retrieve these nodes. We then employ a GNN to aggregate these 2-hop neighborhood signals, as shown in Step (3) in Algotithm 1. Specifically, we first convert textual features of each node (e.g., entity descriptions or relevant text) into dense vector representations using BERT. Let $\mathbf{h}_{1h}$ denote the representation for a 1-hop neighbor $e_{1h}$, while $\mathbf{h}_{2h}$ stands for a 2-hop neighbor $e_{2h}$. We adopt a GraphSAGE-like aggregation hamilton2017inductive to update the representation of each 1-hop neighbor with its 2-hop neighbors:
@@ 5 Experiments
Question-Decomposion huang2023question 72.8 - - -
Chain-of-Question yixing2024chain 78.8 78.10 - -
KG-CoT zhao2024kg 62.3 84.90 68.00 -
FlexKBQA li2024flexkbqa - 46.20 - 68.90
Interactive-KBQA xiong-etal-2024-interactive 49.07 71.2 - -
ToG sun2024thinkongraph 69.5 82.6 57.90 81.4
@@
We evaluate our approach on four standard KBQA datasets, all of which require multi-hop reasoning: ComplexWebQuestions (CWQ) talmor2018web, WebQuestionsSP (WebQSP) yih2016value, WebQuestions berant2013semantic, and GrailQA gu2021beyond. Among them, GrailQA is also used to assess the generalization ability of the retrieval model. All these datasets are based on the Freebase knowledge graph bollacker2008freebase. Following previous studies, we adopt Hit@1 score as the primary evaluation metric.
Implementation Details. For the LLMs, we utilize GPT-4o (gpt-4o-2024-05-13), one of the most advanced general-purpose LLMs, alongside DeepSeek-R1 (70B), a recently released, powerful open-source reasoning model guo2025deepseek. Additionally, we incorporate LLaMA 3.1-70B (hereafter referred to as LLaMA 70B) and a smaller-scale model, LLaMA 3.2-3B (referred to as LLaMA 3B), for comparative analysis.